Define T:R^3 by
T(x,y,z)= (-x,x-y,3x+2y+z).
Check whether T satisfies the polynomial (x-1) (x+1)^2. Find the minimal polynomial of T.
1
Expert's answer
2019-03-22T12:33:55-0400

Denoting by "I" the identity operation, we have

"(T + I) (x, y, z) = (0, x, 3x + 2y + 2z) \, , \\ 
(T + I)^2 (x, y, z) = (0, 0, 8x + 4y + 4z) \, , \\
(T - I) (x, y, z) = (- 2 x, x - 2 y, 3x + 2y) \, ."

Combining these equations together, we obtain

"(T - I) (T + I)^2 (x, y, z) \\ {} = (T - I) (0, 0, 8 x + 4 y + 4 z) = (0, 0, 0) \, ."

Hence, "T" satisfies the polynomial "p (x) = (x - 1) (x + 1)^2". Considering "v = (1, 0, 0)", we have

"T (v) = (-1, 1, 3) \, , \quad T^2 (v) = (1, -2, 2) \, ."

The vectors "v", "T (v)" and "T^2 (v)" are linearly independent; hence, the minimal polynomial of "T" has degree three. Since "p (x)" has degree three and is a monic polynomial (its leading coefficient is equal to 1), and since "p (T) = 0", it is the minimal polynomial.